OpenStudy (anonymous):

Solve the system, using substitution. 3a - b = 11 and 2a + 3b = 0 A. (3, 2) B. (3, -2) C. (-3, 2) D. (-3, -2) E. no solution

OpenStudy (anonymous):

need help with finding the right answer please

OpenStudy (ajspeller):

Assuming (a,b) .. sub in the values for the a and b variables. The results must be true for both equations.

OpenStudy (anonymous):

So in substitution we need to define one of the variables ( a or b) and substitute it in the other equation

OpenStudy (anonymous):

The the first equation 3a-b=11 and change to show b = ????

OpenStudy (anonymous):

Plug them in and you'll get the answer o:

OpenStudy (pinksapphire):

Substitute the multiple answer choices to see if the equations are true.

OpenStudy (anonymous):

Oh feh xD Did not see the second one~

OpenStudy (pinksapphire):

So do the first one by plugging (3,2) in both of the equations to see if they are true and if they are not true then move on to the next equation.

OpenStudy (anonymous):

If I am allowed to..I will explain the first option and you can go from there

OpenStudy (pinksapphire):

I'll start off with the first equation, if you don't know how to do it. 2( )-( )=11 2( )+3( )=0 Plug in (3,2) in for (a,b) so assume that a=3 and b=2 2(3)-(2)=11 2(3)+3(2)=0 6-2=11 6+6=0 4=11 12=0 No No The first set of given coordinates don't work so do the next set of coordinates.

OpenStudy (triciaal):

to use substitution from one equation write the equivalent expression for one of the variables use this expression in the other equation instead of writing the letter of the variable you know have one equation with one unknown variable solve for this variable then substitute this value in the expression to calculate the value of the other variable.

OpenStudy (triciaal):

for this problem from equation 1. b = 3a - 11 use this in equation 2 2a + 3(3a - 11) = 0 11a = 33 so a = 3 b = 3(3) - 11 = 9-11= -2 (3, -2)